# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v2

@@ 5.1 Used ML Models and Key Findings
Table 4 summarizes various research methods or models used in risk prediction and analysis, along with the key findings and associated metrics from different research papers. Each row in the table represents a research paper’s first author, the research method or model used, and the specific findings and metrics achieved. These methods and findings are crucial in understanding risk prediction in various domains.
Table 4: Performance Evaluation of Various Models and Embeddings.
Reference(s) Methods/Models Key Findings
Katsaliaki et al., (2022)
Tensorization-based neural symbolic model
Outperformed SNLP and GNN-LP in predicting relationships (AUC: 0.82, AP: 0.72).
Wang et al., 2022b
Cost-sensitive learning to a random forest (CSL-RF) model
Improved prediction accuracy (Accuracy: 97.22%, AUC: 0.9850).
Wang et al., (2021)
Adaptive heterogeneous multiview graph learning
Improved forecasting accuracy (Accuracy: 87.44%).
Jianying et al., (2021)(Jianying et al., 2021)
Optimized BP neural network
PSO-BP outperformed single BP (MAE: 2.1022, R2: 0.932).
Chen et al., (2021)
LSTM for forecasting
LSTM model suitable for risk prediction (MAE: 1.037 - 1.52).
Zhu et al., (2016)
RS-RAB
Best prediction performance (Accuracy: 86.74%).
Bassiouni et al., (2023)
Various DL approaches
DL models improved accuracy by 14.46% (Softmax: 76.32%, SVM: 85.23%).
Zhu et al., (2019)(Zhu et al., 2019)
RS-MultiBoosting
superior performance of the RS-MultiBoosting ensemble method in forecasting
Athaudage et al., (2022)
Random Forest regression
RF suitable for price prediction (MAE: 1.077, RMSE: 1.56, MAPE: 4.03%).
Liu and Huang, (2020)
AdaBoosted SVM (EN-AdaPSVM)
Outperforms other SVM models in noise data (Total False Rate: 11.76% (EN-AdaPSVM), 13.41% (AdaPSVM), 32.94% (CSVM)).
Zhu et al., (2017)
Ensemble methods (RS-boosting and multi-boosting)
RS-boosting outperforms other methods (Average Accuracy: 85.41% (RS-boosting), 84.08% (multi-boosting), 74.80% (boosting)).
Fu et al., (2022)
Backpropagation NN (BP), Support Vector Machines (SVM), Genetic Algorithm (GA)
BP-GA model outperforms BP and SVM (Accuracy: BP-GA 97.19%, BP 82.33%, SVM 95.31%).
Zhang, (2022)
Multimodal deep learning (DeepRisk)
DeepRisk outperforms baseline methods.
Yang et al., (2021)
Lasso-logistic model
Lasso-logistic model has best prediction accuracy (Prediction Accuracy: 96.5%, Type II Error: 0.037).
Zhang et al., (2022)
Fuzzy Support Vector Machine (FSVM)
FSVM outperforms traditional SVM and DT (Higher accuracy in training and test sets).
Wang, (2021)
Blockchain and fuzzy neural networks
Average Accuracy: 84.11%, 84.66%(Independent Variables Set Vs).
Li, (2022)
Back Propagation Neural Network (BPNN), Logistic Regression
BPNN and logistic regression provide accurate predictions (Prediction Accuracy: 92.8%).
Bodendorf et al., (2023)
Deep learning model with causal inference
High predictive performance and causal analysis.
Liu et al., (2022)(Liu et al., 2022)
Hybrid model (XGBoost–SMOTENC–RF)
Development of a hybrid model chain, incorporating XGBoost, SMOTENC, and Random Forest models (overall accuracy rate 91.67% and the average percentage error is 6.39%).
Rao and Li, (2022)
Logistic regression, DT, integrated Logistic-RS
Integrated RS improves prediction performance (Accuracy of Logistic: 0.831, Accuracy of Logistic-RS: 0.871, Accuracy of DT: 0.772, Accuracy of DT-RS: 0.795).
Handfield et al., (2020)
Newsfeed analysis
Predicted factory risks for a five-year planning horizon.
Pan and Miao, (2023)
Deep learning BP neural network
BPNN has high accuracy and low relative error (The maximum relative error calculated by the AHP is 57.41%, and the minimum is 5.88%, while the maximum relative error obtained by the BPNN is only 0.00210526%).
Wong et al., (2022)
PLS-SEM, ANN
AI-based models (AIRM) are significant determinants of RP and SCA.
Wang and Song, (2022)
Interval Type 2 Fuzzy Neural Network (IT2FNN)
IT2FNN punishes periodic deception, high accuracy (Accuracy 89%).
Nezamoddini et al., (2020)
Genetic algorithm with NN
Improved profits, lower inventory levels.
Xia et al., (2023)
SVM, Random Forest, MLP, Logistic Regression
Random Forest showed best accuracy and low type I error.
Zhao and Li, (2022)
SVM and BP neural network
Effective risk measurement, low relative error.
Li and Fu, (2022)
PCA-GA-SVM
PCA-GA-SVM outperforms SVM and GA-SVM, high accuracy.
Duan et al., (2021)
BP neural network optimized by genetic algorithm
Better fitting effect, higher prediction precision, and higher convergence speed.
Zhang et al., (2021)
SVM optimized by firefly algorithm (FA-SVM)
Improved classification efficiency and reduced error rates.
Luo et al., (2022)
Support vector machine, Particle Swarm Optimization, AdaBoost
Improved classification effect, highest accuracy of 96.13%.
Dang et al., (2022)
Deep learning technology (BP neural network) and blockchain
Effective application of deep learning and blockchain.
Hosseini and Barker, (2016)
Bayesian network (BN)
BN allows disruption and improvement analysis.
Gao et al., (2022)
Autoregressive integrated (AR), Mixture Density Networks (MDN) model, Autoregressive Integrated Moving Average (ARIMA) model, Multilayer Perceptron-Long short term memory (MLP-LSTM) model, Particle Swarm Optimization (PSO) algorithm and the improved PSO (IPSO) algorithm
AR-MDN outperforms other models, IPSO has strong performance.
Wei, (2022)
Machine learning-based linear regression algorithm (ML-LRA)
ML-LRA reduces supplier credit risk.
Zheng et al., (2023)
Federated learning with dimensionality reduction (PCA)
Collective risk prediction benefits organizations with small datasets (Maximum Difference: 0.0670 (FL vs. Local Learning), 0.0634 (FL vs. Centralized Learning).
Deiva Ganesh and Kalpana, (2022)
Text-mining
Use of social media data for contemporary risk understanding.
Zhang et al., (2023)
Time-decayed long short-term memory (TD-LSTM), Deep Neural Network
Improved predictability through TD-LSTM interpolation.
Baryannis et al., 2019a
Support vector machines (SVM) and tree learning
Framework offers good performance with interpretable and black-box methods (SVM (Support Vector Machine) Accuracy: 0.943, DT (Decision Tree) Accuracy: 0.950, RDT (Restricted Decision Tree) Accuracy: 0.916).
Han and Zhang, (2021)
BP neural network model
BP neural network effectively predicted high to extremely high risks in a food production company’s supply chain.
Belhadi et al., 2021b
Hybrid EML approach: Gama Test (GT), Rotation Forest algorithm (RotF), Loogit Boosting algorithm (LB)
Good performance of RotF-LB model (Best results for RotF-LB: Accuracy 91.72, Precision 92.96, Recall 95.53, F-measure 94.23, AUC 92.89).
Wu et al., 2022b
BPNN and GA-BPNN Credit Risk Model
Comparing iterations and risk prediction accuracy (GA-BPNN risk prediction accuracy above 0.92).
Lei et al., (2023)
Binary image difference common area (BCAoID), SVM-SMA
Highlight the effectiveness of the CGOA-SVM-SMA algorithm (Precision: 85.38%, F-score: 63%, TNR: 72%).
Yin et al., (2022)
Convolutional neural network model (CNN)
Optimal accuracy at 200 iterations (Comprehensive accuracy: 94.7%).
Janjua et al., (2023)
Bi-LSTM CRF model
Comparing Bi-LSTM CRF with Baseline CRF (the Bi-LSTM CRF model achieved an overall F1-Score of 85%, the Baseline CRF model achieved an overall F1-Score of 80%).
Yao et al., (2022)
Hybrid model combining FS-MRI, SVME-AIR, and SMV
AUC and KS scores for different parameters (AUC score of 0.8772 when the parameter ”num_AIR” was set to 10, KS score of 0.6363 when the parameter ”num_AIR” was set to 4).
Zhang et al., (2015)
Support vector machine (SVM) and BP neural network
SVM model is superior to the BP neural network model (BP neural network Accuracy 66.67%, SVM model Accuracy 76.67%).
Aboutorab et al., (2022)
RL-PRI approach
Time efficiency of RL-PRI vs. manual process (Tarif Dispute Accuracy: 0.941463, Recession Accuracy: 0.907937, Sea Level Rise Accuracy: 0.964286, Flood Accuracy: 0.856322).
5.2 Clustering of the ML Models
Table 5 presents a clustering of ML models utilized in SCRA, categorizing them into diverse methodological clusters. It offers a comprehensive view of various approaches employed, including ensemble learning, deep learning, time series models, fuzzy logic models, reinforcement learning, dimensionality reduction, genetic algorithms, Bayesian models, and other innovative approaches used in this domain.
Table 5: Clustering of ML Models in SCRA.
Clusters Methods/Models Used Reference(s)
Ensemble Learning
Random Forest (RF)
(Wang et al., 2022b, )
RS-boosting
(Zhu et al.,, 2016)
Multi-boosting
(Zhu et al.,, 2016)
AdaBoosted SVM (EN-AdaPSVM)
(Liu and Huang,, 2020)
XGBoost–SMOTENC–RF
(Liu et al.,, 2022)
Deep Learning
DeepRisk
(Zhang,, 2022)
Deep learning BPNN-based model
(Pan and Miao,, 2023)
Hybrid model (XGBoost–SMOTENC–RF)
(Liu et al.,, 2022)
Time-decayed long short-term memory (TD-LSTM)
(Zhang et al.,, 2023)
Deep Neural Network
(Fu et al.,, 2022)
LSTM
(Chen et al., 2021)
Support Vector Machines
Support vector machine (SVM)
(Zheng et al.,, 2023)
SVM and tree learning
(Baryannis et al., 2019b, )
SVM optimized by firefly algorithm (FA-SVM)
(Zhang et al.,, 2021)
SVM, Particle Swarm Optimization, Adaboost
(Luo et al.,, 2022), (Pan and Miao,, 2023)
Hybrid Models
Hybrid ensemble ML approach: Gama Test (GT), Rotational Forest Algorithm (RotF), Logit Boosting algorithm(LB)
(Belhadi et al., 2021b, )
Blockchain and fuzzy neural networks
(Wang,, 2021)
CGOA-SVM-SMA model for financial risk in the supply chain
(Lei et al.,, 2023)
Time Series Models
AR-MDN
(Gao et al.,, 2022)
ARIMA
(Gao et al.,, 2022)
LSTM for forecasting
(Chen et al.,, 2021), (Gao et al.,, 2022)
Fuzzy Logic Models
Fuzzy Support Vector Machine (FSVM)
(Zhang et al.,, 2022)
Interval Type 2 Fuzzy Neural Network (IT2FNN)
(Wang and Song,, 2022)
Neural Networks
Back Propagation Neural Network (BPNN)
(Li,, 2022), (Zhang et al.,, 2015),(Fu et al.,, 2022)
Backpropagation (BP) NN, SVM, Genetic Algorithm (GA)
(Fu et al.,, 2022)
BP neural network optimized by GA
(Duan et al.,, 2021)
BPNN and logistic regression
(Li,, 2022)
Causal inference in supplier disruption analysis
(Bodendorf et al.,, 2023)
Bayesian network (BN)
(Hosseini and Barker,, 2016)
Text-mining
(Deiva Ganesh and Kalpana,, 2022)
Binary image difference common area (BCAoID)
(Lei et al.,, 2023)
Hybrid model combining FS-MRI, SVME-AIR and SVM
(Yao et al.,, 2022)
Reinforcement Learning
Proactive risk identification using Reinforcement Learning (RL-PRI)
(Aboutorab et al.,, 2022)
Dimensionality Reduction
Federated Learning (FL) with dimensionality reduction (PCA)
(Zheng et al.,, 2023)
Genetic Algorithm
GA with NN
(Nezamoddini et al.,, 2020)
BP NN and GA
(Hosseini and Barker,, 2016)
Bayesian Models
Bayesian network (BN)
(Hosseini and Barker,, 2016)
Other Models
RF regression
(Athaudage et al.,, 2022)
SVM, RF, MLP, Logistic Regression
(Xia et al.,, 2023)
6 Managerial Implications
The findings of this study have significant implications for managerial strategies aimed at enhancing SCRM. Firstly, it underscores the pivotal role of ML and AI in revolutionizing SCRA. Managers are encouraged to explore the adoption of advanced ML models, such as Random Forest, XGBoost, and hybrid approaches, as they have demonstrated substantial potential in improving risk assessment precision. These models can significantly bolster the accuracy of risk evaluation and enhance the overall effectiveness of risk management strategies.
Moreover, the research emphasizes the importance of adaptability and flexibility in the post-COVID era, highlighting the need for proactive supply chain strategies. Managers should prioritize the development of contingency plans that are resilient to potential future disruptions, ensuring the continuity of operations and minimizing the impact of unforeseen events.
The study also underscores the effectiveness of ensemble models, such as Random Forest and XGBoost, along with hybrid approaches in risk assessment. These models can potentially enhance risk prediction precision and should be considered by managers seeking to fortify their SCRM. Time series forecasting techniques, including ARIMA and LSTM, emerge as pivotal in anticipating future risks. Organizations are encouraged to invest in predictive models grounded in historical data, providing them with the foresight required to mitigate supply chain disruptions effectively.
The significance of XAI and ML models cannot be overstated. Ensuring that models offer clear and interpretable results is essential, as this aids decision-makers in comprehending the risk assessment process and its outcomes. The study also emphasizes the importance of real-time data mining and incorporating unconventional data sources, such as social media platforms like Twitter, as essential elements of comprehensive risk identification. Managers must stay vigilant, continually updating their understanding of real-time information.
Blockchain technology emerges as a potent tool for enhancing transparency and traceability within the supply chain. Organizations considering blockchain solutions should evaluate their applications for risk mitigation, ultimately enhancing SCR. SMEs warrant tailored strategies for credit risk assessment, accounting for their unique operational challenges. Implementing strategies aligned with the specific needs and constraints of SMEs is vital. In cases where data privacy is a concern, the adoption of federated learning is recommended. This approach safeguards sensitive information while still enabling collaborative risk prediction. Ensuring data security and privacy compliance should be a top priority.
Hybrid risk assessment methodologies, which integrate various ML models and techniques, offer a holistic view of SCRA. Managers should explore these comprehensive approaches to build robust risk management strategies. Benchmarking different ML models is crucial for identifying the most effective ones for specific risk scenarios. The study underscores the significance of continuous performance evaluation and improvement, ensuring that risk assessment strategies evolve with changing circumstances. Integrating ML-based risk assessment into existing supply chain workflows and decision-making processes is pivotal to ensure seamless implementation. Managers should explore ways to align these technologies with their current operational strategies. Recognizing the importance of employee training in data analysis, ML, and AI is fundamental. Managers should invest in skill development to empower their supply chain professionals to harness these technologies effectively. Collaboration and knowledge sharing within organizations and across industries are encouraged to collectively build robust risk management strategies. Such partnerships have the potential to enhance risk resilience significantly.
Lastly, understanding and adhering to data protection and privacy regulations is critical when deploying AI and ML for risk assessment. Managers should ensure that their risk management practices are fully aligned with the legal requirements to avoid any compliance issues. The study offers a comprehensive roadmap for organizations aiming to fortify their SCRM strategies in an ever-evolving landscape.

### v3

@@ 6.1 Used ML Models and Key Findings
Table LABEL:tab:ml_methods summarizes various research methods or models used in risk prediction and analysis, along with the key findings and associated metrics from different research papers. Each row in the table represents a research paper’s first author, the research method or model used, and the specific findings and metrics achieved. These methods and findings are crucial in understanding risk prediction in various domains.
6.2 AI Models Used in SCRA
The application of AI in SCRA has seen significant growth, driven by the need for robust, predictive models to manage risk in complex and dynamic supply networks. To effectively analyze the methodologies employed in the reviewed literature, we classified the AI models into five main groups: CML Models, DL Models, Hybrid Models, Ensemble Methods, and Advanced Ensemble Methods. Each category is defined based on its distinct methodological characteristics, level of complexity, application scope, and suitability for specific SC risk scenarios, as shown in Figure 15.
Figure 15: Classification of AI models in SCRA based on methodology, complexity, and common application scenarios.
A detailed overview of each AI model used in SCRA, including their key characteristics, advantages, disadvantages, application cases, and suitable scenarios, is given below:
6.2.1 Conventional ML (CML) Models
Support Vector Machine (SVM): SVM is a supervised ML algorithm effective in high-dimensional spaces. It finds a hyperplane to separate classes in the feature space, often using kernel functions for non-linear separation. SVM performs well with limited data, making it useful for SCs with restricted historical data. By maximizing the margin, SVMs are less prone to overfitting, especially in smaller datasets. However, SVM can be computationally expensive for large datasets, which can be limiting in SCs with extensive data. Its results are also difficult to interpret, especially with complex kernel functions, which can be challenging for explainable predictions in SC risk. Bassiouni et al., (2023) found that the SVM classifier achieved an accuracy of 84.56% in predicting SC risks. Fu et al., (2022) found that the SVM model achieved a classification accuracy of 95.31% in assessing credit risk in Internet-based SC financing, outperforming the BPNN but slightly underperforming the BP-GA model. SVM is ideal for high-stakes, high-dimensional risk scenarios, like financial risk assessment with limited historical data. The data type is structured, often with limited samples. Input variables include supplier attributes, transaction history, and financial indicators, with the target variable being binary risk classification (e.g., high-risk vs. low-risk suppliers).
Decision Tree (DT): DTs are a tree-based algorithm that splits data into branches based on feature values, resulting in a flowchart-like model that is easy to interpret. DTs offer a simple structure that is easy to interpret, making it valuable in SC risk scenarios that require transparency. They are also relatively fast to train and execute, which makes them suitable for real-time or near-real-time assessments. However, DT models are prone to overfitting, especially with noisy data, and may require pruning or ensemble methods to mitigate this. They also struggle with capturing complex, non-linear relationships compared to models like neural networks. Rao and Li, (2022) found that the DT model achieved an accuracy of 77.2% in predicting SC financing risks. However, while highly interpretable, Baryannis et al., 2019a found that DT learning achieved lower performance predicting SC risks than SVM. The DT model demonstrated a trade-off between interpretability and performance, with SVM providing superior predictive accuracy, particularly in binary classification tasks for late deliveries. DT models are adequate for straightforward, interpretable assessments of SC risk, especially with structured and smaller datasets. The data type is structured with categorical or binary risk indicators. Input variables include financial ratios, transaction amounts, and supplier location, with the target variable being risk level classification (e.g., high, medium, low).
Logistic Regression (LR): LR is a linear model that estimates probabilities for categorical outcomes, commonly used for binary classification in risk assessment. LR offers a straightforward, easy-to-interpret linear approach that is beneficial for transparent decision-making in SC risk. It also requires less data and computational power, making it suitable for smaller datasets. However, LR is unsuited for complex, non-linear relationships, which can reduce accuracy in intricate SC risk scenarios. It also assumes feature independence, which might not hold in interconnected SC networks. Li, (2022) found that the LR model achieved a high prediction accuracy of 92.8% in forecasting investment risks in SCM, demonstrating excellent applicability and potential for broader use in SCM risk prediction. While Yang et al., (2021) found that the Logistic Regression model achieved a prediction accuracy of 83.1% in evaluating the credit risk of SMEs in SC finance, the lasso logistic model significantly outperformed it with an accuracy of 96.5% and a lower Type II error rate. LR is ideal for binary risk scenarios and cases where model transparency is essential. The data type is structured, typically tabular data. Input variables include supplier payment history, financial indicators, and supply volume, with the target variable being binary risk classification (e.g., default vs. no default).
6.2.2 Deep Learning (DL) Models
DL models, like neural networks, use multiple layers to model complex, non-linear relationships, making them practical for complex SC risk predictions. DL models are ideal for intricate, non-linear relationships in SC data, such as demand forecasting under uncertain conditions. They generally achieve high accuracy in well-established scenarios, particularly with large datasets. However, the “black-box” nature of DL models can make them difficult to interpret, which may limit adoption in risk-sensitive contexts. They also require large datasets and computational power, hindering some SC applications. Pan and Miao, (2023) found that the BPNN model demonstrated excellent performance in SCRA, achieving a maximum relative error of only 0.0021%, significantly outperforming the AHP model, which had a maximum relative error of 57.41%. This indicates the BPNN’s superior ability to evaluate risk accurately and reliably in SC systems. DL is suitable for large-scale, complex SCRAs with high data availability and less need for interpretability. The data type is large, structured, or unstructured data. Input variables include time series data on SC metrics, inventory levels, and demand trends, with the target variable being the probability of disruption or a risk score.
6.2.3 Hybrid Models (CML + DL Models)
Hybrid models combine CML and DL techniques to leverage the strengths of each for more robust SCRA. Hybrid models often outperform single models by combining strengths, such as using DL for feature extraction and CM for interpretability. They are versatile across various scenarios by tailoring components to specific data characteristics and risk types. However, hybrid models can be challenging to implement and tune, requiring expertise in multiple algorithms. Additionally, combining models can increase computational costs, which may be challenging for real-time applications. Fu et al., (2022) found that the hybrid BPNN and GA model outperformed both the standalone BPNN and SVM models in predicting credit risk for internet-based SC financing. The BPNN-SVM hybrid achieved the highest classification accuracy of 97.19%, significantly improving prediction performance compared to the BPNN (82.33%) and SVM (95.31%) models. Hybrid models are well-suited for applications requiring both high accuracy and moderate interpretability, especially when handling complex, multi-source data. The data type is mixed, including structured and semi-structured data. Input variables include supplier reliability, lead times, transactional volumes, and customer demand, with the target variable being multi-class risk categorization (e.g., high, medium, low).
6.2.4 Ensemble Methods
Ensemble methods like Bagging, Boosting, and Stacking combine multiple models to improve accuracy and robustness in SC risk prediction. Ensemble methods reduce variance and bias by aggregating multiple models, which makes them suitable for high-stakes risk assessment. They are versatile, as they can work with various algorithms, allowing for tailored approaches depending on the data and risk scenario. However, ensemble methods can require substantial computational resources, particularly for Boosting and Stacking. While they boost accuracy, they can decrease interpretability, making it difficult to understand individual model contributions. Liu and Huang, (2020) utilized the EN-AdaPSVM model, integrating AdaBoost with SVM and enhanced by noise elimination and adaptive mutation particle swarm optimization, achieving high classification accuracy and robustness in SC finance risk assessment. Luo et al., (2022) highlighted that while their AdaBoost-DPSO-SVM model showed significant performance improvements, it faced challenges such as higher error rates in negative sample recognition, suggesting potential issues with bias and misclassification. Ensemble methods are ideal for high-accuracy, high-stakes SC risk scenarios, especially when model interpretability is not the primary concern. The data type is structured, tabular data, often large datasets. Input variables include financial history, supplier performance scores, and geographic risk indicators, with the target variable being binary or multi-class risk levels (e.g., high-risk, low-risk, moderate risk).
6.2.5 Advanced Ensemble Methods
XGBoost: XGBoost is a powerful gradient-boosting algorithm known for its speed and accuracy, especially in structured data applications. XGBoost often outperforms traditional algorithms in structured data tasks, providing robust risk predictions. It also handles missing data well, which is useful in SC datasets that may be incomplete. However, XGBoost models can be complex, reducing interpretability and making tuning more challenging. Liu et al., (2022) developed a credit risk assessment model for the construction industry’s online SC finance using a hybrid approach combining XGBoost, SMOTENC, and RF. The study aimed to address the gaps in existing credit assessment systems, which often rely on traditional models with low accuracy and poor interpretability. Using data from 71 upstream enterprises between 2018 and 2020, the model achieved an overall accuracy rate of 91.67% on test samples. XGBoost is effective for structured, high-dimensional datasets in SC risk, particularly where interpretability is secondary to accuracy. The data type is large, structured data, often tabular. Input variables include financial indicators, transactional history, and geographic data, with the target variable being risk classification or score.
Random Forest (RF): RF is an ensemble method based on DTs, known for reducing variance and improving predictive accuracy. RF retains interpretability compared to more complex models and effectively handles high-dimensional data while being less prone to overfitting. However, RF can become slow with large datasets, particularly when building large forests. High-depth trees can also lead to overfitting if not carefully tuned. Wang et al., (2022) utilized a cost-sensitive Random Forest (CSL-RF) model to predict SMEs’ credit risk in SC finance, incorporating 40 diverse predictors and employing an imbalance sampling strategy for high accuracy. Athaudage et al., (2022) applied RF regression to forecast crude oil prices during disease outbreaks by integrating disease data with oil market variables, demonstrating reliable performance under complex conditions. Both studies highlight the effectiveness of RF models in handling varied data sets for robust forecasting. RF is effective when interpretability is moderately important and high accuracy and robustness are needed. The data type is structured, often large-scale. Input variables include supplier attributes, financial ratios, and location-based risk factors, with the target variable being multi-class risk level (e.g., high, medium, low).
7 Practical and Managerial Implications
The findings of this article will play a significant role in managerial strategies designed to enhance SCRM. First, it stresses that ML and AI are important tools that will give rise to a new phase of SCRA. Managers are encouraged to try advanced ML models, such as RF, XGBoost, and blends of these. They stimulate considerable risk assessment precision. These models are likely to enhance the accuracy of risk evaluation enormously and, thus, the effectiveness of their risk management strategies.
In addition, the study highlights the importance of flexibility and adaptability in an era after COVID-19, suggesting that companies need new SC strategies. Managers must now place contingency plans at the forefront, focusing on those not readily broken by any future disruptions. It is only by doing routine work thoroughly that one can concentrate on something special without fear or disturbance from the outside world, using people who have already developed a variety of fresh paths, so avoiding crisis after crisis with all that effort having been wasted.
The findings also highlight the importance of ensemble methods – RF and XGBoost – and hybrid methodologies in risk assessment. It is possible that such models could improve the accuracy of risk prediction and may be of interest to managers striving to strengthen their SCRM systems. The importance of time series prediction models, such as ARIMA and LSTM, cannot be overstated. Companies are advised to create tools predicting potential risks based on historical performance to approach the task of SC disruption mitigation with much-needed foresight.
XAI and ML models are extremely important. Models must produce precise and human-readable outputs as they assist decision-makers in understanding the process of risk assessment and even its findings. Furthermore, the study indicates the significance of real-time data mining and using non-traditional sources like social media, e.g., Twitter, as critical components of full-scale risk identification. Managers should be careful and up-to-date in understanding real-time information.
Blockchain represents an effective instrument for increasing transparency and traceability in the SC, thereby making risk monitoring mechanisms more efficient. Organizations exploring blockchain solutions should determine the possibilities of addressing and using the risk to boost SCR. SMEs require principal strategies in risk evaluation in terms of CR since they have distinct characteristics and risks in terms of operations. Conversely, implementing strategies based on SMEs’ needs and constraints is required. Moreover, federated learning should be implemented when data privacy issues arise. This method addresses the risk but protects confidential information, permitting partners to conduct joint risk forecasting. Compliance with security and data privacy must be performed almost immediately.
Several hybrid risk assessment methodologies that combine various ML models and techniques bring the necessary comprehensive view of SCRA, offering managers the chance to look into such approaches to develop risk management strategies. Additionally, comparing benchmarks of different ML models is critical to understanding more information about the best ones for particular risk scenarios. The overall message of the study is outlining the need for continuous performance evaluations and improvements. Risk assessment strategies should be able to evolve alongside other circumstances. Other assumptions refer to integrating ML-based risk assessment into already existing SC workflows and decision-making patterns. It is necessary to determine how such technologies can be easily implemented. Employee training in data analysis, ML, and AI is critical. Managers should invest in skill development to empower their SC professionals to harness these technologies effectively. Collaboration and knowledge sharing within organizations and across industries are encouraged to collectively build robust risk management strategies. Such partnerships have the potential to enhance risk resilience significantly.
Lastly, understanding and adhering to data protection and privacy regulations is critical when deploying AI and ML for risk assessment. Managers should ensure that their risk management practices align with the legal requirements to avoid compliance issues. The study offers a comprehensive roadmap for organizations aiming to fortify their SCRM strategies in an ever-evolving landscape.